OpenStudy (mimi_x3):

Can you please help me with this maths question 1. The cubic polynomial P(x)= x^3 +rs^2 +sx+t , where r,s and t are real numbers has three real zeroes,1,a, and -a (a) find the value of s + t

myininaya (myininaya):

is the polynomial written correctly?

OpenStudy (mimi_x3):

woops my bad sorry P(x) = x^3 +rx^2 +sx +t

OpenStudy (helpme1234):

Colorful pi can you help me?

myininaya (myininaya):

let me try to help him and then i will look at your question

OpenStudy (mimi_x3):

him ? are yu referring to me ?

myininaya (myininaya):

is this algebra?

OpenStudy (mimi_x3):

noo , its a polynomial question

myininaya (myininaya):

so what class is it then?

OpenStudy (mimi_x3):

class ? what do yu mean ? Its from the topic polynomials

myininaya (myininaya):

\[P(x)=(x-1)(x-a)(x+a)=x^3+rs^2+sx+t\]

OpenStudy (mimi_x3):

is that the answer ?

myininaya (myininaya):

so I guess you are doing it for fun? \[P(x)=(x-1)(x^2-a^2)=x^3+rs^2+sx+t\] no i just thought u would like to see the steps

OpenStudy (mimi_x3):

no , im not doing it for fun , its my maths assignment LOL

myininaya (myininaya):

\[P(x)=x(x^2-a^2)-1(x^2-a^2)=x^3+rsx^2+sx+t\]

myininaya (myininaya):

so the class is called polynomials lol?

OpenStudy (mimi_x3):

what does yu mean by class ?

myininaya (myininaya):

\[P(x)=x^3-a^2x-x^2+a^2=x^3+rsx^2+sx+t\] people go to class to learn and get assignments lol

OpenStudy (mimi_x3):

no the class is called maths LOL

myininaya (myininaya):

\[P(x)=x^3-x^2-a^2x+a^2=x^3+rsx^2+sx+t\]

OpenStudy (mimi_x3):

2 unit maths to be specific

myininaya (myininaya):

ok we are almost done with this problem can you guess what to do next?

myininaya (myininaya):

we want both sides to be the same

OpenStudy (mimi_x3):

collecting like terms ?

myininaya (myininaya):

ok the the coefficient of x^2 on left hand side is -1 the coefficient of x^2 on the other sides is rs so what can you say about rs and -1?

myininaya (myininaya):

they must be equal

myininaya (myininaya):

so we need that for the other terms as well

myininaya (myininaya):

we have -1=rs -a^2=s a^2=t

OpenStudy (anonymous):

I seriously think I'm being trolled here.. but just to ease my tired brain.. \[\sqrt{25} \ne -5\]

myininaya (myininaya):

you are right polpak

OpenStudy (mimi_x3):

wait , dnt yu factorise it further to (x^2-a^2)(x-1)

myininaya (myininaya):

that =5

myininaya (myininaya):

factor?

OpenStudy (anonymous):

Someone was trying to tell me that it equals both 5 and -5

myininaya (myininaya):

ok the objective is to find s+t

OpenStudy (anonymous):

But I can only really prove it by going to the definition. Which they don't accept.

OpenStudy (mimi_x3):

ohh yeah , my bad

OpenStudy (anonymous):

Sorry for bothering you mimi, but myin rarely watches group chat (with good reason)

myininaya (myininaya):

what do you think s+t is based on what i gave you above? hint we don't need that first equation rs=-1

OpenStudy (mimi_x3):

yu talking to me or polpak ? LOL

myininaya (myininaya):

you

OpenStudy (mimi_x3):

lols , tbh im lost , i dont get it xDD

myininaya (myininaya):

do you understand when I wrote P as P(x)=(x-1)(x-a)(x+a)?

OpenStudy (mimi_x3):

yeah , i dont understand the part -1 = rs -a^2 = s a^2 = t where did yu get it from ?

myininaya (myininaya):

both sides had to be equal we were looking at \[P(x)=x^3-a^2x-x^2+a^2=x^3+rsx^2+sx+t \]

myininaya (myininaya):

look at the coefficients of each term on both sides we want each term to be the same as it is on the other side

myininaya (myininaya):

so if we look at the x^3's they are good they both have coefficient 1 what is the coefficient of x^2 on left hand side ? what is the coefficient of x^2 on right hand side?

OpenStudy (mimi_x3):

ohh i get it rs = -1

myininaya (myininaya):

yes now what is the coefficient of x on the left hand side? what is the coefficient of x on the right hand side?

OpenStudy (mimi_x3):

x on the left hand side: 1 i think x on the right hand side 1 i think

myininaya (myininaya):

no what is the number in front of x on boths?

myininaya (myininaya):

both sides*

OpenStudy (mimi_x3):

there's no other number

myininaya (myininaya):

-a^2? s?

OpenStudy (mimi_x3):

lols , its a letter not a number

myininaya (myininaya):

they are numbers

OpenStudy (mimi_x3):

they are ? arent they are called pronumerals ?

OpenStudy (mimi_x3):

are they*

myininaya (myininaya):

never heard of this term

OpenStudy (mimi_x3):

ohh ok

myininaya (myininaya):

when you are doing math and you don't see a word and you see a letter it usually represents a number unless it is the word a or i

myininaya (myininaya):

becauce a and i can be words

OpenStudy (mimi_x3):

ohh ok , i never knew that they are numbers xD

myininaya (myininaya):

but usually the book will make it clear to you what they are

myininaya (myininaya):

yep

OpenStudy (mimi_x3):

ok , so x= a^3

myininaya (myininaya):

a^3 is nowhere to be found we aren;t solving equation for x

OpenStudy (mimi_x3):

ok , so x= a^3

OpenStudy (mimi_x3):

ohh ok

myininaya (myininaya):

ok so look again at the polynomial what is the coefficent of x on both sides?

OpenStudy (mimi_x3):

ok , so x= a^3

OpenStudy (mimi_x3):

ok , so x= a^3

OpenStudy (mimi_x3):

x= s

myininaya (myininaya):

ok no \[P(x)=x^3-a^2x-x^2+a^2=x^3+rsx^2+sx+t \] ^ ^ | | -a^2 = s

OpenStudy (mimi_x3):

ok , so x= a^3

OpenStudy (mimi_x3):

ohh lols , sorry im so slow of understanding it

myininaya (myininaya):

ok do you see the terms that don't an x in em? there terms are called constants and they must be equal of that we have the same polynomial on both sides

myininaya (myininaya):

god my english is getting bad

OpenStudy (mimi_x3):

ok , so x= a^3

OpenStudy (mimi_x3):

lols , probably getting annoyed of me not getting it xD

myininaya (myininaya):

no its just late here i been up for along time today

myininaya (myininaya):

a long*

OpenStudy (anonymous):

Ditto!

myininaya (myininaya):

so what are the constants on both sides?

myininaya (myininaya):

the terms that don't have a x in them

OpenStudy (mimi_x3):

ok , so x= a^3

OpenStudy (mimi_x3):

t

OpenStudy (mimi_x3):

ok , so x= a^3

myininaya (myininaya):

t and what is it on the other side?

myininaya (myininaya):

remember you are looking for non-x terms

OpenStudy (mimi_x3):

ok , so x= a^3

OpenStudy (mimi_x3):

a

myininaya (myininaya):

a or a^2?

OpenStudy (mimi_x3):

a

myininaya (myininaya):

\[P(x)=x^3-a^2x-x^2+a^2=x^3+rsx^2+sx+t \] ^ ^ | | a^2 = t

OpenStudy (mimi_x3):

a^2 my bad

myininaya (myininaya):

so what is s again? and what is t again?

OpenStudy (mimi_x3):

s = 1 t = a^2

myininaya (myininaya):

s is not 1